A 2.0 kg ball and a 3.0 kg ball are connected by a 1.5 m long rigid, massless rod. The rod and balls are rotating clockwise about its center of gravity at 16 rpm . Part A What torque will bring the balls to a halt in 9.0 s ?
Answer #1

Center of mass, r = (2 x 0 + 3 x 1.5)/(2 + 3) = 0.9 m

Distance from center of rod, d = 1.5 - 0.9 = 0.6 m

Moment of inertia, I = (2 x 0.9^2) + (3 x 0.6^2) = 2.7 kg m^2

Angular speed, w = 16 x 2 pi/60 = 1.68 rad/sec

Angular acceleration, a = 1.68/9 = 0.186 rad/s^2

Torque, T = 2.7 x 0.186 = 0.50 Nm

Comment in case any doubt please rate my answer ..